OpenStudy (anonymous):

Derivative of (1+x)^(1/x)

OpenStudy (anonymous):

have you done something called the D-operator?

OpenStudy (anonymous):

first of all rewrite that as (1+x)^(-x), does that clue or hint help?

OpenStudy (anonymous):

in other words you may use implicit differentiation ...

OpenStudy (anonymous):

so then it's chain rule

OpenStudy (anonymous):

ok just for you to never forget do you know how to find dy/dx of x^x?

OpenStudy (anonymous):

yes using natural logs

OpenStudy (anonymous):

ok then use the same method and differnetiate both sides

OpenStudy (anonymous):

so it would be -1-2x

OpenStudy (anonymous):

impossible to give that answer life is not linear...

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

y = (1+x)^(-x)

OpenStudy (anonymous):

take the ln of both sides...

OpenStudy (anonymous):

ln(y) = ln[(1+x)^(-x)]

OpenStudy (anonymous):

my appology that the equation editor is not functioning for me tonight for a funny reason...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

use natural logs rules to obtain: ln(y)=-xln(1+x)

OpenStudy (anonymous):

now can you diff both sides implicitly ...

OpenStudy (anonymous):

f'(x)=-ln(1+x)-[x/(1+x)]*(1+x)^(-x)

OpenStudy (anonymous):

no look at what i wrote again ln(y)=-xln(1+x)

OpenStudy (anonymous):

you used the product rule but u made a small error

OpenStudy (anonymous):

let's work on this step by step

OpenStudy (anonymous):

ok sure

OpenStudy (anonymous):

now the editor is working

OpenStudy (anonymous):

\[ \ln(y)=-xln(1+x)\]

OpenStudy (anonymous):

\[d( \ln(y)=-xln(1+x))\]

OpenStudy (anonymous):

(1/y)y'=-1*ln(x+1)=(-x)(1/(1+x))

OpenStudy (anonymous):

you mean + not =

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok all you are left with is y x RHS

OpenStudy (anonymous):

take the y to the other side

OpenStudy (anonymous):

y'=y[-1*ln(x+1)=(-x)(1/(1+x))]

OpenStudy (anonymous):

can you see the y?

OpenStudy (anonymous):

yes... everything** is multiplied by y, not just tacked onto the end

OpenStudy (anonymous):

the next question is what is y=?

OpenStudy (anonymous):

just change the y to the main equation

OpenStudy (anonymous):

now I have to apply that to x -> 0, L'Hopital's Rule

OpenStudy (anonymous):

no

OpenStudy (anonymous):

delete that post

OpenStudy (anonymous):

we don't have limits

OpenStudy (anonymous):

\[y'=[(1+x)^(1/x)][-1*\ln(x+1)=(-x)(1/(1+x))]\]

OpenStudy (anonymous):

can u see i changed the y to (1+x)^(1/x)

OpenStudy (anonymous):

yes, i got that

OpenStudy (anonymous):

good let's us wrap it up

OpenStudy (anonymous):

in such questions 1) you take the natrual log of both side, 2) you apply the logs rules, 3) you implicitly differentiate both sides, 4) you plug in the oridinal equation or function back

OpenStudy (anonymous):

one of the interesting things about this phenomina, that you'll have the same function muliplied by it derivative in the logarithmic world, a lot of real life applications can be modeled by this system...

OpenStudy (anonymous):

Ok, great... now the whole point of that was to answer the question: lim x->0 (1+0)^(1/x), and applying l'hopital's rule

OpenStudy (anonymous):

that is a different question now

OpenStudy (anonymous):

the answer would be 1/e

OpenStudy (anonymous):

it is easy to see

OpenStudy (anonymous):

Please show me how it is easy to see.

OpenStudy (anonymous):

sorry the answer would be e

OpenStudy (anonymous):

now one way to visualise this is by graphical method and the definition of limits, the other way to find out is by using calculus normal methods

OpenStudy (anonymous):

ok to make life esear

OpenStudy (anonymous):

you should use LH rule

OpenStudy (anonymous):

let me prove that for you

OpenStudy (anonymous):

yes, I need to understand this.

OpenStudy (anonymous):

\[y = (1+x)^{(1/x)}\]

OpenStudy (anonymous):

\[\ln(y) =\ln[ (1+x)^{(1/x)}]\]

OpenStudy (anonymous):

\[\ln(y) =\frac{\ln[ (1+x)}{x}\]

OpenStudy (anonymous):

i actually just took a quiz on this

OpenStudy (anonymous):

are you at hopkins by chance? haha... I want to understand this

OpenStudy (anonymous):

\[\ln(y) = \lim_{x \rightarrow 0} [\frac{\ln[ (1+x)}{x}]\]

OpenStudy (anonymous):

i say (1+x)^1/x-1/x

OpenStudy (anonymous):

im prob wrong tho lol

OpenStudy (anonymous):

now differentiate the top and the butoom using LH rule...

OpenStudy (anonymous):

can you see what will happen

OpenStudy (anonymous):

(1/1+x)/x

OpenStudy (anonymous):

\[\ln(y) = \lim_{x \rightarrow 0} [\frac{\frac{1}{[ (1+x)}}{1}]\]

OpenStudy (anonymous):

yes plug in the 0

OpenStudy (anonymous):

oh yes, 1

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

ln y =1

OpenStudy (anonymous):

y =e

OpenStudy (anonymous):

hehehehe

OpenStudy (anonymous):

I see!

OpenStudy (anonymous):

its a clever qs

OpenStudy (anonymous):

this is the first time is see such a problem, but just follow the laws of maths, the it will drive us mad and crazy ...

OpenStudy (anonymous):

then it will*

OpenStudy (anonymous):

y = e^1

OpenStudy (anonymous):

which means y =e

OpenStudy (anonymous):

did you get that...or am i disconnected...

OpenStudy (anonymous):

no, i see that... much more advanced that what I'm used to seeing, but it makes sense.

OpenStudy (anonymous):

well good luck and all the best

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

where is my medal hehehe